OpenStudy (cherrilyn):

√xy=x^2+1.. what is dy/dx? , @calculus

OpenStudy (anonymous):

\[\sqrt{xy} =x ^{2}+1\]

OpenStudy (anonymous):

find the implicit differentiation

OpenStudy (cherrilyn):

yeah I square rooted both sides and did that... I got dy/dx= 4x^2 - y/x but I think it's wrong

OpenStudy (cherrilyn):

i mean squared****

myininaya (myininaya):

\[\sqrt{xy}=x^2+1\] \[\frac{1}{2}(xy)^\frac{-1}{2}(xy)'=2x+0\] \[\frac{1}{2 \sqrt{xy}}(y+xy')=2x\]

myininaya (myininaya):

\[\frac{y}{2 \sqrt{xy}}+\frac{xy'}{2 \sqrt{xy}}=2x\]

myininaya (myininaya):

\[\frac{xy'}{2 \sqrt{xy}}=2x-\frac{y}{ 2 \sqrt{xy}}\]

myininaya (myininaya):

\[y'=\frac{2 \sqrt{xy}}{x}(2x-\frac{y}{2 \sqrt{xy}})\]

myininaya (myininaya):

\[y'=4 \sqrt{xy}-\frac{y}{x}\]

myininaya (myininaya):

\[y'=\frac{4x \sqrt{xy}-y}{x}\]

myininaya (myininaya):

now let's try your way squaring both sides \[xy=(x^2+1)^2\] now differentiate both sides \[y+xy'=2(x^2+1)(2x+0)\] \[y+xy'=4x(x^2+1)\] \[y+xy'=4x \sqrt{xy}\] \[xy'=4x \sqrt{xy} -y\] \[y'=\frac{4 x \sqrt{xy} -y}{x}\] samething :)

myininaya (myininaya):

\[\text{ remember } \sqrt{xy}=x^2+1\]

myininaya (myininaya):

so i replaced x^2+1 with (xy)^(1/2) to show we would get the same result

myininaya (myininaya):

hey money did you delete what you have did you have \[y+xy'=4x(x^2+1)\] \[y+xy'=4x^3+4x\] \[xy'=4x^3+4x-y\] \[y'=\frac{4x^3+4x-y}{x}=4x^2+4-\frac{y}{x}\] this is not wrong

OpenStudy (cherrilyn):

not right you mean?

myininaya (myininaya):

this is not wrong means this is right

myininaya (myininaya):

this is one of the ways to write the y'

OpenStudy (cherrilyn):

my book says that to prove that the equation (dy/dx) = \[4\sqrt{x} - (y/x)\]

myininaya (myininaya):

i have that above

OpenStudy (cherrilyn):

so I'm confused

myininaya (myininaya):

just divide each term by x

myininaya (myininaya):

i wrote this above \[y'=\frac{4 x \sqrt{xy} -y}{x}=4 \sqrt{xy} -\frac{y}{x}\]

OpenStudy (cherrilyn):

but there's an extra y

myininaya (myininaya):

you mean \[4 \sqrt{xy}-\frac{y}{x} \text{ \right?}\]

myininaya (myininaya):

i went off money's equation is that equation written incorrectly?

OpenStudy (cherrilyn):

oh , FAIL I overlooked it. YES That's exactly it! thank you so much myininaya :D